# Entropic vector quantile regression: Duality and Gaussian case

### v1

@@ 4. Proofs for Section 2
$P(A)\leq\frac{a+\log 2}{\log(1/Q(A))}.$
Now, since $Q$ is $\beta$-sub-Weibull, there exists a finite constant $K^{\prime}$ such that
$Q(\{x:\|x\|\geq t\})\leq K^{\prime}e^{-t^{\beta}/K^{\prime}},\ t>0.$
For sufficiently large $t$,
$\frac{1}{\log(1/Q(A))}$
decays like $t^{-\beta}$. Combining this estimate with
$\begin{split}\int\|x\|\mathbbm{1}_{\{\|x\|\geq r\}}\,dP(x)&=\int_{M}\int_{0}^{\infty}\mathbbm{1}_{\{\|x\|\geq r\vee t\}}\,dtdP(x)\\ &=rP(\{x:\|x\|\geq r\})+\int_{r}^{\infty}P(\{x:\|x\|\geq t\})\,dt\end{split}$
one can conclude that $P_{n}$ is uniformly integrable. ∎
We are now in position to prove Lemma 2.1.
@@
Case (i) directly follows from Lemma 4.1 combined with the fact that $e^{-c/\varepsilon}$ is bounded. So we focus on Case (ii). From Lemma 4.1, it suffices to verify that $\tilde{R}$ is $\beta$-sub-Weibull. Recall that $d\tilde{R}(u,x,y)=\alpha^{-1}e^{-c(u,y)/\varepsilon}\,d\mu(u)d\nu(x,y)$ and observe that $c(u,y)\geq-\|u\|^{2}/2+\|y\|^{2}/4$. Let $U\sim\mu$ be independent of $(X,Y)$. Then for any $K>0$,
$\mathbb{E}\big[e^{K(\|U\|^{\beta}+\|X\|^{\beta}+\|Y\|^{\beta})}e^{-c(U,Y)/\varepsilon}]\big]\leq\mathbb{E}\big[e^{K\|U\|^{\beta}+\|U\|^{2}/(2\varepsilon)}\big]\mathbb{E}\big[e^{K\|X\|^{\beta}}e^{K\|Y\|^{\beta}-\|Y\|^{2}/(4\varepsilon)}\big].$
Since $U$ is compactly supported, $\mathbb{E}\big[e^{K\|U\|^{\beta}+\|U\|^{2}/(2\varepsilon)}\big]<\infty$. Furthermore, for $K<1/(4\varepsilon)$, $e^{K\|y\|^{\beta}-\|y\|^{2}/(4\varepsilon)}$ is bounded, so that the right-hand side on the displayed equation is bounded, up to a constant, by $\mathbb{E}[e^{K\|X\|^{\beta}}]$, which is finite for a sufficiently small $K$ as $X$ is $\beta$-sub-Weibull. This implies that $\tilde{R}$ is $\beta$-sub-Weibull. ∎
4.4. Proof of Theorem 2.1
@@
We shall translate the feasible set into countably many unconditional moment constraints. To this end, we first verify the following lemma. Recall the feasible set $\mathcal{Q}$ in (7).
@@
One can find countably many functions $\{(\mathfrak{f}_{i},\mathfrak{g}_{i},\mathfrak{h}_{i}):i\in\mathbb{N}\}$ with $\mathfrak{f}_{i}:\mathbb{R}^{d_{y}}\to[-1,1],\mathfrak{g}_{i}:\mathbb{R}^{d_{y}}\to[-1,1]^{d_{x}},\mathfrak{h}_{i}:\mathbb{R}^{d_{x}+d_{y}}\to[-1,1]$ being $1$-Lipschitz, such that the following holds: for a given $\pi\in\mathcal{P}_{1}(\mathbb{R}^{d_{y}}\times\mathbb{R}^{d_{x}+d_{y}})$,
@@
The first and third constraints in (22) ensure that $\pi\in\Pi(\mu,\nu)$. Constructing such functions $\mathfrak{f}_{i}$ and $\mathfrak{h}_{i}$ is standard. For example, recalling that the collection $\mathsf{BL}_{1}(\mathbb{R}^{d_{y}})$ of $1$-Lipschitz functions $\mathbb{R}^{d_{y}}\to[-1,1]$ is separable for the topology of uniform convergence on compacta by the Ascoli-Arzelà theorem, we take $\{\mathfrak{f}_{i}:i\in\mathbb{N}\}$ to be its countable dense subset. Then the first constraint in (22) holds for all $i\in\mathbb{N}$ if and only if the first marginal of $\pi$ agrees with $\mu$. The construction of $\mathfrak{h}_{i}$ is similar.
@@
$\begin{split}\int_{\mathbb{R}^{d_{x}+d_{y}}}x\,d\pi_{u}(x,y)=0\ \text{$\mu$-a.e. $u$}&\Longleftrightarrow\int_{\mathbb{R}^{d_{y}}\times\mathbb{R}^{d_{x}+d_{y}}}\mathbbm{1}_{A}(u)x\,d\pi(u,x,y)=0,\ \forall A\in\mathcal{B}(\mathbb{R}^{d_{y}})\\ &\Longleftrightarrow\int_{\mathbb{R}^{d_{y}}\times\mathbb{R}^{d_{x}+d_{y}}}f(u)x\,d\pi(u,x,y)=0,\ \forall f\in\mathsf{BL}_{1}(\mathbb{R}^{d_{y}}).\end{split}$
@@
$\begin{split}&\int_{\mathbb{R}^{d_{y}}\times\mathbb{R}^{d_{x}+d_{y}}}f(u)x\,d\pi(u,x,y)=0,\ \forall f\in\mathsf{BL}_{1}(\mathbb{R}^{d_{y}})\\ &\Longleftrightarrow\int_{\mathbb{R}^{d_{y}}\times\mathbb{R}^{d_{x}+d_{y}}}\langle\mathfrak{g}_{i,j}(u),x\rangle\,d\pi(u,x,y)=0,\ \forall i\in\mathbb{N},\forall j\in\{1,\dots,d_{x}\}.\end{split}$
@@
By [CSI75, Theorem 3.1], there exist $(f_{n},g_{n},h_{n})\in L^{\infty}(\mu)\times L^{\infty}(\mu;\mathbb{R}^{d_{x}})\times L^{\infty}(\nu)$ such that
@@
$F(u,x,y):=\log\frac{d\pi^{\varepsilon}}{d\tilde{R}}(u,x,y)=\lim_{n\to\infty}\big(f_{n}(u)+\langle g_{n}(u),x\rangle+h_{n}(x,y)\big)\in[-\infty,\infty)$
@@
$(u,x,y)\in S_{0}\Longrightarrow(u,x_{*},y_{*})\in S_{0}\ \text{and}\ (u_{*},x,y)\in S_{0}.$
Now, if we define
$\begin{split}&\tilde{f}_{n}(u)=f_{n}(u)-f_{n}(u_{*}),\\ &\tilde{g}_{n}(u)=g_{n}(u)-g_{n}(u_{*}),\\ &\tilde{h}_{n}(x,y)=h_{n}(x,y)+f_{n}(u_{*})+\langle g_{n}(u_{*}),x\rangle,\end{split}$
then $\tilde{f}_{n}(u_{*})=0,\tilde{g}_{n}(u_{*})=0$, and
@@
For notational convenience, relabel $(\tilde{f}_{n},\tilde{g}_{n},\tilde{h}_{n})$ as $(f_{n},g_{n},h_{n})$.
Set $\mathcal{U}_{00}:=\{u\in\mathcal{U}_{0}:(u,x,y)\in S_{0}\ \text{for some}\ (x,y)\in\mathcal{Z}_{0}\}$ and $\mathcal{Z}_{00}:=\{(x,y)\in\mathcal{Z}_{0}:(u,x,y)\in S_{0}\ \text{for some}\ u\in\mathcal{U}_{0}\}$. These sets need not be Borel measurable, but are analytic and hence $\mu$- and $\nu$-completion measurable, respectively (cf. Chapter 13 in [DUD02]). If necessary, remove null sets from $\mathcal{U}_{00}$ and $\mathcal{Z}_{00}$ so that they are Borel. Observe that $\mu(\mathcal{U}_{00})=\nu(\mathcal{Z}_{00})=1$. Pick any $(x,y)\in\mathcal{Z}_{00}$. Since $(u_{*},x,y)\in S_{0}$, we have
$F(u_{*},x,y)=\lim_{n\to\infty}F_{n}(u_{*},x,y)=\lim_{n\to\infty}h_{n}(x,y).$
@@
$h(x,y):=\begin{cases}F(u_{*},x,y)&\text{if $(x,y)\in\mathcal{Z}_{00}$,}\\ 0&\text{otherwise},\end{cases}$
@@
Lemma 4.3.
One can find $d_{x}+1$ points $(x_{0},y_{0}),\dots,(x_{d_{x}},y_{d_{x}})\in\mathcal{Z}_{00}$ such that the vectors $(1,x_{0}^{\top})^{\top},\dots,(1,x_{d_{x}}^{\top})^{\top}$ form a basis of $\mathbb{R}^{d_{x}+1}$.
Proof of Lemma 4.3.
If there are no such $d_{x}+1$ vectors, then the distribution of $(1,X^{\top})^{\top}$ is concentrated on an $d_{x}$-dimensional linear subspace of $\mathbb{R}^{d_{x}+1}$. As such, there exists a constant nonzero vector $w=(w_{1},w_{2}^{\top})^{\top}\in\mathbb{R}^{d_{x}+1}$ such that $w_{1}+w_{2}^{\top}X=0$ a.s. Observe that $w_{2}$ cannot be zero; otherwise $w_{1}=0$, contradicting $w\neq 0$. This, however, implies that $X$ is concentrated on an affine hyperplane, contradicting Assumption 2.1 (i). ∎
Let $V$ be the invertible $(d_{x}+1)\times(d_{x}+1)$ matrix whose rows are $v(x_{i},y_{i})^{\top}$, and define
$\varphi_{n}(u):=\begin{pmatrix}\langle H_{n}(u),v(x_{0},y_{0})\rangle\\ \vdots\\ \langle H_{n}(u),v(x_{d_{x}},y_{d_{x}})\rangle\end{pmatrix}=VH_{n}(u).$
@@
$\big(f(u),g(u)^{\top}\big)^{\top}:=\lim_{n\to\infty}H_{n}(u)=\lim_{n\to\infty}V^{-1}g_{n}(u)=V^{-1}\varphi(u),\ u\in\mathcal{U}_{00},$
and set $(f(u),g(u)^{\top})^{\top}=0$ outside $\mathcal{U}_{00}$. We have shown that
@@
Now, by Lemma 4.3, one can find $d_{x}+1$ points $x_{0},x_{1},\dots,x_{d_{x}+1}$ such that $K(x_{i})$ are finite for all $i\in\{0,\dots,d_{x}\}$, and that $v_{i}:=x_{i}-x_{0}$ ($i\in\{1,\dots,d_{x}\}$) form a basis of $\mathbb{R}^{d_{x}}$. Since $K(x_{i})-K(x_{0})=\int_{\mathbb{R}^{d_{y}}}\langle g(u),v_{i}\rangle\,d\mu(u)$, we see that $\langle g(\cdot),v_{i}\rangle\in L^{1}(\mu)$ for all $i\in\{1,\dots,d_{x}\}$. This implies that $g\in L^{1}(\mu;\mathbb{R}^{d_{x}})$ and $f\in L^{1}(\mu)$. Finally, since $|h(x,y)|\leq|\tilde{F}(x,y)|+\|f\|_{L^{1}(\mu)}+\|g\|_{L^{1}(\mu)}\|x\|$, $\tilde{F}\in L^{1}(\nu)$, and $\mathbb{E}[X^{2}]<\infty$, we conclude that $h\in L^{1}(\nu)$. In addition, since $|\langle g(u),x\rangle|\leq|F(u,x,y)|+|f(u)|+|h(x,y)|$ and $F\in L^{1}(\pi^{\varepsilon})$ by (24), we have that $\langle g(u),x\rangle$ is integrable under $\pi^{\varepsilon}$ .

### v2

@@ 4. Proofs for Section 2
$P(A)\leq\frac{a+\log 2}{\log(1/Q(A))},$
which implies that $\mathcal{L}_{a}$ is uniformly tight and hence weakly precompact by Prohorov’s theorem.
To show that $\mathcal{L}_{a}$ is compact in $\mathsf{W}_{1}$, it remains to establish that
$\lim_{\lambda\to\infty}\sup_{P\in\mathcal{L}_{a}}\int d(\cdot,x_{0})\mathbbm{1}_{\{d(\cdot,x_{0})>\lambda\}}\,dP=0.$ (24)
See Theorem 6.9 in [VIL09]. Set $\varphi_{\lambda}:=d(\cdot,x_{0})\mathbbm{1}_{\{d(\cdot,x_{0})>\lambda\}}$. By the weighted Pinsker inequality from Theorem 2.1 in [BV05], we have
$\begin{split}\int\varphi_{\lambda}\,dP&\leq\int\varphi_{\lambda}\,dQ+\frac{2}{\alpha}\left(\frac{3}{2}+\log\int e^{\alpha\varphi_{\lambda}}\,dQ\right)\left(\sqrt{\mathsf{KL}(P\,\|\,Q)}+\frac{1}{2}\mathsf{KL}(P\,\|\,Q)\right)\\ &\leq\int\varphi_{\lambda}\,dQ+\frac{2}{\alpha}\left(\frac{3}{2}+\log\int e^{\alpha\varphi_{\lambda}}\,dQ\right)\left(\sqrt{a}+\frac{a}{2}\right),\quad\forall P\in\mathcal{L}_{a}.\end{split}$
Since $d(\cdot,x_{0})\in L^{1}(Q)$ by Condition (23), the dominated convergence theorem yields
$\lim_{\lambda\to\infty}\int\varphi_{\lambda}\,dQ=0.$
In addition,
$\int e^{\alpha\varphi_{\lambda}}\,dQ=Q(\{d(\cdot,x_{0})\leq\lambda\})+\int_{\{d(\cdot,x_{0})>\lambda\}}e^{\alpha d(\cdot,x_{0})}\,dQ.$
Again, by the dominated convergence theorem, the right-hand side converges to 1 as $\lambda\to\infty$. By taking the supremum over $P\in\mathcal{L}_{a}$ and taking the limits $\lambda\to\infty$ and $\alpha\to\infty$ in this order, we obtain (24). ∎
Remark 4.2 (Necessity of Condition 12 in Lemma 2.1).
Assumption 2.1 (ii) may hold even when Condition 12 is not met. To see this, observe that, for $\alpha^{\prime}>0$,
$\int e^{\alpha^{\prime}\tilde{\mathsf{d}}(\cdot,0)}\,d\tilde{R}\leq\alpha^{-1}e^{2\alpha^{\prime}}\int e^{\alpha^{\prime}\|x\|-c(u,y)/\varepsilon}\,d\mu(u)d\nu(x,y).$
Since $c(u,y)\geq-\|u\|^{2}/2+\|y\|^{2}/4$, we have
$\int e^{\alpha^{\prime}\|x\|-c(u,y)/\varepsilon}\,d\mu(u)d\nu(x,y)\leq\underbrace{\int e^{\|u\|^{2}/(2\varepsilon)}\,d\mu(u)}_{=I}\times\underbrace{\int e^{\alpha^{\prime}\|x\|-\|y\|^{2}/(2\varepsilon)}\,d\nu(x,y)}_{=II}.$
The first term II is finite if, e.g., $\mu$ is compactly supported. The second term IIII is finite for any $\alpha^{\prime}>0$ if, e.g., $Y=X$, even when Condition (12) fails to hold.
4.4. Proof of Theorem 2.1
4.4.1. Preliminary lemmas
Before the proof, we shall recall the following (standard) result. We provide its proof for completeness.
Lemma 4.2.
Let $M$ be a metric space and $f:M\to(-\infty,\infty]$ be coercive in the sense that the sublevel sets of $f$ are compact. If $\inf f>$-$\infty$, then there exists at least one $\bar{x}\in M$ such that $f(\bar{x})=\inf f$.
@@ 4.4.1. Preliminary lemmas
The conclusion is trivial if $f\equiv\infty$, so we assume that $a:=\inf f$ is finite. Let $x_{n}$ be a sequence such that $f(x_{n})\to a$. Since $x_{n}\in\{f\leq a+1\}$ for sufficiently large $n$, there exists a convergent subsequence $x_{n_{k}}$, $x_{n_{k}}\to\bar{x}$. Coercivity (in our definition) implies lower semicontinuity, so
[…]
completing the proof. ∎
We will also use the following lemma, which is a small modification to (a special case of) Theorem 3.1 in [CSI75].
Lemma 4.3.
Let $\Omega$ be a measurable space and $\mathcal{P}(\Omega)$ be the set of all probability measures on $\Omega$. For given measurable functions $\phi_{i}:\Omega\to\mathbb{R}\ (i\in\{1,\dots,n\})$, let $\mathcal{Q}\subset\mathcal{P}(\Omega)$ be a nonempty convex set such that $\int\phi_{i}\,dP=0$ for all $i\in\{1,\dots,n\}$ and for all $P\in\mathcal{Q}$. For a given $R\in\mathcal{P}(\Omega)$, suppose that there exists a (unique) $Q\in\mathcal{Q}$ such that $\mathsf{KL}(Q\,\|\,R)=\inf_{P\in\mathcal{Q}}\mathsf{KL}(P\,\|\,R)<\infty$, and that
$\mathcal{Q}^{\prime}:=\left\{P\in\mathcal{P}(\Omega):P\ll Q,\frac{dP}{dQ}\leq 2,\int\phi_{i}\,dP=0,\forall i\in\{1,\dots,n\}\right\}\subset\mathcal{Q}.$ (25)
Then $Q$ has a density of the form
$\frac{dQ}{dR}=ae^{\sum_{i=1}^{n}b_{i}\phi_{i}}$
for some $a>0$ and $b_{i}\in\mathbb{R}$.
Proof of Lemma 4.3.
Theorem 3.1 in [CSI75] assumes that $\mathcal{Q}$ (in our context) is the set of all $P\in\mathcal{P}(\Omega)$ such that $\int f_{i}\,dP=0$ for all $i$. In our case, we allow $\mathcal{Q}$ to be smaller than that, but Condition (25) ensures that $Q$ minimizes $\mathsf{KL}(\cdot\,\|\,R)$ over $\mathcal{Q}^{\prime}$, which contains $Q$ as an algebraic inner point. So, mimicking the argument in the proof of Theorem 3.1 in [CSI75] yields the claim of the lemma (recall that any finite-dimensional subspace of $L^{1}(Q)$ is closed in $L^{1}(Q)$). ∎
The following lemma is inspired by Lemma 2.10 in [NUT21]. In the proof of Theorem 2.1 below, we will construct dual potentials from the $R$-a.e. limit of $f_{n}(u)+\langle g_{n}(u),x\rangle+h_{n}(x,y)$ for suitable measurable functions $(f_{n},g_{n},h_{n})$. Because of the presence of the interaction term $\langle g_{n}(u),x\rangle$, special care is needed to construct dual potentials in such a way that they are separately measurable, i.e., one has to choose “canonical” points more carefully than standard entropic OT.
Lemma 4.4.
Suppose Assumption 2.1 (i) holds. If a Borel subset $S\subset\mathbb{R}^{d_{y}}\times\mathbb{R}^{d_{x}+d_{y}}$ has full $R$-measure, then one can choose $u^{*}\in\mathbb{R}^{d_{y}},(x_{i}^{*},y_{i}^{*})\in\mathbb{R}^{d_{x}+d_{y}}\ (i\in\{0,1,\dots,d_{x}\})$ with $(u^{*},x_{i}^{*},y_{i}^{*})\in S$ for all $i\in\{0,1,\dots,d_{x}\}$ for which the following hold:
  1. (i)
    the $d_{x}+1$ vectors
    $\binom{1}{x_{0}^{*}},\dots,\binom{1}{x_{d_{x}}^{*}}$
    form a basis of $\mathbb{R}^{d_{x}+1}$; and
  2. (ii)
    there exist $\mathcal{U}_{0}\subset\mathbb{R}^{d_{y}}$ and $\mathcal{Z}_{0}\subset\mathbb{R}^{d_{x}+d_{y}}$ with $\mu(\mathcal{U}_{0})=\nu(\mathcal{Z}_{0})=1$ such that, for $S_{0}:=(\mathcal{U}_{0}\times\mathcal{Z}_{0})\cap S$, it holds that $(u^{*},x_{i}^{*},y_{i}^{*})\in S_{0}$ for all $i\in\{0,1,\dots,d_{x}\}$ and
    $(u,x,y)\in S_{0}\Longrightarrow(u^{*},x,y)\in S_{0}\ \text{and}\ (u,x_{i}^{*},y_{i}^{*})\in S_{0},\ \forall i\in\{0,1,\dots,d_{x}\}.$
Proof of Lemma 4.4.
For notational convenience, we write $z=(x,y)$. For $u\in\mathbb{R}^{d_{y}}$, let $S_{u}:=\{z:(u,z)\in S\}$. Define $S_{z}$ analogously. By Fubini’s theorem, $\nu(S_{u})=1$ for $\mu$-a.e. $u$. Pick any $u^{*}$ for which $\nu(S_{u^{*}})=1$, and set $\mathcal{Z}_{0}:=\{z:\mu(S_{z})=1\}\cap S_{u^{*}}$, which has full $\nu$-measure. By Lemma 4.5 below, one can choose $z_{0}^{*},\dots,z_{d_{x}}^{*}\in\mathcal{Z}_{0}$ (with $z_{i}^{*}=(x_{i}^{*},y_{i}^{*}))$ for which (i) holds. Now, set $\mathcal{U}_{0}:=\{u:\nu(S_{u})=1\}\cap S_{z_{0}^{*}}\cap\cdots\cap S_{z_{d_{x}}^{*}}$, which has full $\mu$-measure. By construction, the sets $\mathcal{U}_{0},\mathcal{Z}_{0}$ satisfy (ii). ∎
Lemma 4.5.
Suppose Assumption 2.1 (i) holds. If a Borel subset $\mathcal{Z}_{0}\subset\mathbb{R}^{d_{x}+d_{y}}$ has full $\nu$-measure, then one can find $d_{x}+1$ points $(x_{0},y_{0}),\dots,(x_{d_{x}},y_{d_{x}})\in\mathcal{Z}_{0}$ such that the vectors $(1,x_{0}^{\top})^{\top},\dots,(1,x_{d_{x}}^{\top})^{\top}$ form a basis of $\mathbb{R}^{d_{x}+1}$.
Proof of Lemma 4.5.
Let $\mathcal{X}_{0}:=\{x:(x,y)\in\mathcal{Z}_{0}\ \text{for some}\ y\}$. Observe that $\mathcal{X}_{0}$ has full measure with respect to the completion of the law of $X$. The claim of the lemma states that one can find $x_{0},\dots,x_{d_{x}}\in\mathcal{X}_{0}$ such that $(1,x_{0}^{\top})^{\top},\dots,(1,x_{d_{x}}^{\top})^{\top}$ form a basis of $\mathbb{R}^{d_{x}+1}$. Suppose on the contrary that there are no such $d_{x}+1$ vectors. Then $\{(1,x^{\top})^{\top}:x\in\mathcal{X}_{0}\}$ is contained in a $d_{x}$-dimensional vector subspace of $\mathbb{R}^{d_{x}+1}$. Hence, there exists a nonzero vector $w=(w_{1},w_{2}^{\top})^{\top}\in\mathbb{R}^{d_{x}+1}$ such that $w_{1}+w_{2}^{\top}x=0$ for all $x\in\mathcal{X}_{0}$. Observe that $w_{2}$ cannot be zero; otherwise $w_{1}=0$, contradicting $w\neq 0$. This, however, implies that $X$ is concentrated on an affine hyperplane, contradicting Assumption 2.1 (i). ∎
We are now ready to prove Theorem 2.1.
4.4.2. Proof of Theorem 2.1
@@
We shall translate the feasible set into countably many unconditional moment constraints. To this end, we first verify the following lemma. Recall the definitions of $\mathcal{Q}$ in (8) and $\tilde{\mathcal{P}}_{1}$ in (11). For a metric space $M$, let $BL_{1}(M)$ denote the set of all $1$-Lipschitz functions $M\to[-1,1]$. Observe that, when $M$ is $\sigma$-compact, $BL_{1}(M)$ is separable for the topology of uniform convergence on compacta.
@@
One can find countably many functions $\{(\mathfrak{f}_{i},\mathfrak{g}_{i},\mathfrak{h}_{i}):i\in\mathbb{N}\}$, with $\mathfrak{f}_{i}:\mathbb{R}^{d_{y}}\to\mathbb{R},\mathfrak{g}_{i}:\mathbb{R}^{d_{y}}\to\mathbb{R}^{d_{x}}$, and $\mathfrak{h}_{i}:\mathbb{R}^{d_{x}+d_{y}}\to\mathbb{R}$, such that $\mathfrak{f}_{i}\in BL_{1}(\mathbb{R}^{d_{y}}),\mathfrak{h}_{i}\in BL_{1}(\mathbb{R}^{d_{x}+d_{y}})$, and each coordinate of $\mathfrak{g}_{i}$ belongs to $BL_{1}(\mathbb{R}^{d_{y}})$, for all $i\in\mathbb{N}$, and that for a given $\pi\in\tilde{\mathcal{P}}_{1}$,
@@
The first and third constraints in (27) ensure that $\pi\in\Pi(\mu,\nu)$. Constructing such functions $\mathfrak{f}_{i}$ and $\mathfrak{h}_{i}$ is standard. For example, take $\{\mathfrak{f}_{i}:i\in\mathbb{N}\}$ to be a countable dense subset for $BL_{1}(\mathbb{R}^{d_{y}})$ with respect to the topology of uniform convergence on compacta. Then the first constraint in (27) holds for all $i\in\mathbb{N}$ if and only if the first marginal of $\pi$ agrees with $\mu$. The construction of $\mathfrak{h}_{i}$ is similar.
@@
$\begin{split}\int_{\mathbb{R}^{d_{x}+d_{y}}}x\,d\pi_{u}(x,y)=0\ \text{$\mu$-a.e. $u$}&\Longleftrightarrow\int_{\mathbb{R}^{d_{y}}\times\mathbb{R}^{d_{x}+d_{y}}}\mathbbm{1}_{A}(u)x\,d\pi(u,x,y)=0,\ \forall A\in\mathcal{B}(\mathbb{R}^{d_{y}})\\ &\Longleftrightarrow\int_{\mathbb{R}^{d_{y}}\times\mathbb{R}^{d_{x}+d_{y}}}f(u)x\,d\pi(u,x,y)=0,\ \forall f\in BL_{1}(\mathbb{R}^{d_{y}}).\end{split}$
The first equivalence follows by the definition of the conditional expectation. To verify the second equivalence, define a (vector-valued) signed measure on $\mathcal{B}(\mathbb{R}^{d_{y}})$ by $\rho(A):=\int_{\mathbb{R}^{d_{y}}\times\mathbb{R}^{d_{x}+d_{y}}}\mathbbm{1}_{A}(u)x\,d\pi(u,x,y)$. The $\Rightarrow$”direction is trivial. For the $\Leftarrow$” direction, for any closed $A\subset\mathbb{R}^{d_{y}}$, one can approximate $\mathbbm{1}_{A}$ by bounded Lipschitz functions to obtain $\rho(A)=0$. Applying Dynkin’s $\pi$-$\lambda$ theorem, we conclude $\rho(A)=0$ for all $A\in\mathcal{B}(\mathbb{R}^{d_{y}})$.
@@
$\begin{split}&\int_{\mathbb{R}^{d_{y}}\times\mathbb{R}^{d_{x}+d_{y}}}f(u)x\,d\pi(u,x,y)=0,\ \forall f\in BL_{1}(\mathbb{R}^{d_{y}})\\ &\Longleftrightarrow\int_{\mathbb{R}^{d_{y}}\times\mathbb{R}^{d_{x}+d_{y}}}\langle\mathfrak{g}_{i,j}(u),x\rangle\,d\pi(u,x,y)=0,\ \forall i\in\mathbb{N},\forall j\in\{1,\dots,d_{x}\}.\end{split}$
@@
As Condition (25) above holds for $\mathcal{Q}=\mathcal{Q}_{n},R=\tilde{R}$, $Q=\pi_{n}$, and $\phi_{i}$’s suitably chosen, we can apply Lemma 4.3 to conclude that there exist $(f_{n},g_{n},h_{n})\in L^{\infty}(\mu)\times L^{\infty}(\mu;\mathbb{R}^{d_{x}})\times L^{\infty}(\nu)$ such that
@@
These are constructed as follows: $f_{n}$ is a linear combination of $\{1,\mathfrak{f}_{1},\dots,\mathfrak{f}_{n}\}$, and $g_{n},h_{n}$ are linear combinations of $\{\mathfrak{g}_{i}:i\leq n\}$ and $\{\mathfrak{h}_{i}:i\leq n\}$, respectively.
@@
$F(u,x,y):=\log\frac{d\pi^{\varepsilon}}{d\tilde{R}}(u,x,y)=\lim_{n\to\infty}\big(f_{n}(u)+\langle g_{n}(u),x\rangle+h_{n}(x,y)\big)\in[-\infty,\infty)$ (30)
@@
We invoke Lemma 4.4 above to find sets $\mathcal{U}_{0}\subset\mathbb{R}^{d_{y}},\mathcal{Z}_{0}\subset\mathbb{R}^{d_{x}+d_{y}}$ with $\mu(\mathcal{U}_{0})=\nu(\mathcal{Z}_{0})=1$ and points $u^{*}\in\mathbb{R}^{d_{y}},(x_{i}^{*},y_{i}^{*})\in\mathbb{R}^{d_{x}+d_{y}}\ (i\in\{0,1,\dots,d_{x}\})$ for which (i) and (ii) in Lemma 4.4 hold. Recall $S_{0}=(\mathcal{U}_{0}\times\mathcal{Z}_{0})\cap S$. If we define
$\begin{split}&\tilde{f}_{n}(u)=f_{n}(u)-f_{n}(u^{*}),\\ &\tilde{g}_{n}(u)=g_{n}(u)-g_{n}(u^{*}),\\ &\tilde{h}_{n}(x,y)=h_{n}(x,y)+f_{n}(u^{*})+\langle g_{n}(u^{*}),x\rangle,\end{split}$
then $\tilde{f}_{n}(u^{*})=0,\tilde{g}_{n}(u^{*})=0$, and
@@
For notational convenience, relabel $(\tilde{f}_{n},\tilde{g}_{n},\tilde{h}_{n})$ as $(f_{n},g_{n},h_{n})$. Set $\mathcal{U}_{00}:=\{u\in\mathcal{U}_{0}:(u,x,y)\in S_{0}\ \text{for some}\ (x,y)\in\mathcal{Z}_{0}\}$ and $\mathcal{Z}_{00}:=\{(x,y)\in\mathcal{Z}_{0}:(u,x,y)\in S_{0}\ \text{for some}\ u\in\mathcal{U}_{0}\}$. These sets need not be Borel measurable, but are analytic and hence $\mu$- and $\nu$-completion measurable, respectively (cf. Chapter 13 in [DUD02]). Observe that $\bar{\mu}(\mathcal{U}_{00})=\bar{\nu}(\mathcal{Z}_{00})=1$, with $\bar{\mu},\bar{\nu}$ denoting the completions of $\mu,\nu$, respectively.
Pick any $(x,y)\in\mathcal{Z}_{00}$. Since $(u^{*},x,y)\in S_{0}$, we have
$F(u^{*},x,y)=\lim_{n\to\infty}F_{n}(u^{*},x,y)=\lim_{n\to\infty}h_{n}(x,y).$
@@
$h(x,y):=\begin{cases}F(u^{*},x,y)&\text{if $(x,y)\in\mathcal{Z}_{00}$,}\\ 0&\text{otherwise},\end{cases}$
@@
Pick any $u\in\mathcal{U}_{00}$. Observe that $(u,x_{i}^{*},y_{i}^{*})\in S_{0}$ for all $i\in\{0,1,\dots,d_{x}\}$. Let $V$ be the invertible $(d_{x}+1)\times(d_{x}+1)$ matrix whose rows are $v(x_{i}^{*},y_{i}^{*})^{\top}$, and define
$\varphi_{n}(u):=\begin{pmatrix}\langle H_{n}(u),v(x_{0}^{*},y_{0}^{*})\rangle\\ \vdots\\ \langle H_{n}(u),v(x_{d_{x}}^{*},y_{d_{x}}^{*})\rangle\end{pmatrix}=VH_{n}(u).$
@@
$\big(f(u),g(u)^{\top}\big)^{\top}:=\lim_{n\to\infty}H_{n}(u)=\lim_{n\to\infty}V^{-1}\varphi_{n}(u)=V^{-1}\varphi(u),\ u\in\mathcal{U}_{00},$
[…]
@@
Now, by Lemma 4.5, one can find $d_{x}+1$ points $x_{0},x_{1},\dots,x_{d_{x}}$ such that $K(x_{i})$ are finite for all $i\in\{0,\dots,d_{x}\}$, and that $v_{i}:=x_{i}-x_{0}$ ($i\in\{1,\dots,d_{x}\}$) form a basis of $\mathbb{R}^{d_{x}}$. Since $K(x_{i})-K(x_{0})=\int_{\mathbb{R}^{d_{y}}}\langle g(u),v_{i}\rangle\,d\mu(u)$, we see that $\langle g(\cdot),v_{i}\rangle\in L^{1}(\mu)$ for all $i\in\{1,\dots,d_{x}\}$. This implies that $g\in L^{1}(\mu;\mathbb{R}^{d_{x}})$ and $f\in L^{1}(\mu)$. Finally, since $|h(x,y)|\leq|\tilde{F}(x,y)|+\|f\|_{L^{1}(\mu)}+\|g\|_{L^{1}(\mu)}\|x\|$, $\tilde{F}\in L^{1}(\nu)$, and $\mathbb{E}[\|X\|]<\infty$, we conclude that $h\in L^{1}(\nu)$. In addition, since $|\langle g(u),x\rangle|\leq|F(u,x,y)|+|f(u)|+|h(x,y)|$ and $F\in L^{1}(\pi^{\varepsilon})$ by (29), we have that $\langle g(u),x\rangle$ is integrable under $\pi^{\varepsilon}$.